# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v1

@@ H.1 The role of dual attention
Table 9: Directed Graph Generation performance across different transformer architectures using discrete flow matching.
@@
RRWP - Dual 0.0107 $\pm$ 0.0019 0.0104 $\pm$ 0.0012 0.1209 $\pm$ 0.0194 0.0054 $\pm$ 0.0005 0.0043 $\pm$ 0.0008 1.7 $\pm$ 0.1 94 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 94 $\pm$ 1.0
MagLap 0.0113 $\pm$ 0.0024 0.0110 $\pm$ 0.0015 0.1196 $\pm$ 0.0107 0.0061 $\pm$ 0.0011 0.0029 $\pm$ 0.0006 1.7 $\pm$ 0.2 74 $\pm$ 4.4 100 $\pm$ 0.0 100 $\pm$ 0.0 74 $\pm$ 4.4
MagLap - Dual 0.0110 $\pm$ 0.0017 0.0116 $\pm$ 0.0021 0.0509 $\pm$ 0.0017 0.0062 $\pm$ 0.0014 0.0027 $\pm$ 0.0005 1.3 $\pm$ 0.2 91 $\pm$ 2.5 100 $\pm$ 0.0 100 $\pm$ 0.0 91 $\pm$ 2.5
SBM No PE 0.0038 $\pm$ 0.0021 0.0040 $\pm$ 0.0020 0.1912 $\pm$ 0.0178 0.0975 $\pm$ 0.0028 0.0961 $\pm$ 0.0071 20.5 $\pm$ 2.0 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0 $\pm$ 0.0
@@
MagLap - Dual 0.0039 $\pm$ 0.0011 0.0036 $\pm$ 0.0010 0.0653 $\pm$ 0.0026 0.0033 $\pm$ 0.0003 0.0038 $\pm$ 0.0010 1.9 $\pm$ 0.3 77 $\pm$ 7.6 100 $\pm$ 0.0 100 $\pm$ 0.0 77 $\pm$ 7.6
In particular, we observe that, across both datasets, the inclusion of the dual attention mechanism consistently improves the validity of the generated graphs and reduces structural discrepancies as measured by clustering, spectral, and wavelet distances. In the ER-DAG setting, models with dual attention (e.g., RRWP-Dual and MagLap-Dual) significantly outperform their non-dual counterparts in validity—achieving up to 94% validity with RRWP-Dual and 91% with MagLap-Dual, compared to just 67.5% and 74%, respectively. These models also demonstrate better structural fidelity, particularly in clustering and spectral metrics. A similar trend is observed in the SBM dataset, where dual attention boosts validity (e.g., RRWP-Dual: 87%, MagLap-Dual: 77%). These results demonstrate how incorporating bidirectional information flow into the attention mechanism contributes to improved model performance and provides insight into the effectiveness of this architectural component.
Directo-DD
@@
Table 10: Directed Graph Generation performance across different transformer architectures using discrete diffusion.
@@
ER-DAG No PE 0.0094 $\pm$ 0.0024 0.0095 $\pm$ 0.0023 0.1221 $\pm$ 0.0122 0.0814 $\pm$ 0.0030 0.1094 $\pm$ 0.0066 14.4 $\pm$ 0.8 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0 $\pm$ 0.0
No PE - Dual 0.0149 $\pm$ 0.0037 0.0138 $\pm$ 0.0035 0.1256 $\pm$ 0.0088 0.0089 $\pm$ 0.0018 0.0106 $\pm$ 0.0014 2.6 $\pm$ 0.3 53 $\pm$ 2.9 100 $\pm$ 0.0 100 $\pm$ 0.0 53 $\pm$ 2.9
RRWP 0.0147 $\pm$ 0.0038 0.0139 $\pm$ 0.0038 0.0985 $\pm$ 0.0154 0.0067 $\pm$ 0.0009 0.0038 $\pm$ 0.0009 2.1 $\pm$ 0.3 42 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 42 $\pm$ 1.0
RRWP - Dual 0.0142 $\pm$ 0.0040 0.0129 $\pm$ 0.0034 0.0408 $\pm$ 0.0053 0.0055 $\pm$ 0.0010 0.0048 $\pm$ 0.0014 1.4 $\pm$ 0.3 79 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 79 $\pm$ 3.7
MagLap 0.0144 $\pm$ 0.0040 0.0135 $\pm$ 0.0038 0.0792 $\pm$ 0.0126 0.0071 $\pm$ 0.0012 0.0036 $\pm$ 0.0011 1.6 $\pm$ 0.3 59 $\pm$ 6.8 100 $\pm$ 0.0 100 $\pm$ 0.0 59 $\pm$ 6.8
MagLap - Dual 0.0142 $\pm$ 0.0044 0.0135 $\pm$ 0.0034 0.0438 $\pm$ 0.0071 0.0057 $\pm$ 0.0013 0.0046 $\pm$ 0.0010 1.4 $\pm$ 0.3 69 $\pm$ 6.4 100 $\pm$ 0.0 100 $\pm$ 0.0 69 $\pm$ 6.4
SBM No PE 0.0038 $\pm$ 0.0021 0.0040 $\pm$ 0.0020 0.1912 $\pm$ 0.0177 0.0120 $\pm$ 0.0019 0.0975 $\pm$ 0.0081 20.5 $\pm$ 2.0 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0 $\pm$ 0.0
No PE - Dual 0.0037 $\pm$ 0.0020 0.0039 $\pm$ 0.0019 0.0721 $\pm$ 0.0059 0.0055 $\pm$ 0.0007 0.0403 $\pm$ 0.0067 8.8 $\pm$ 1.3 35 $\pm$ 9.1 100 $\pm$ 0.0 100 $\pm$ 0.0 35 $\pm$ 9.1
RRWP 0.0038 $\pm$ 0.0019 0.0038 $\pm$ 0.0018 0.1085 $\pm$ 0.0205 0.0088 $\pm$ 0.0009 0.0359 $\pm$ 0.0043 8.5 $\pm$ 0.3 4 $\pm$ 2.0 100 $\pm$ 0.0 100 $\pm$ 0.0 4 $\pm$ 2.0
@@
H.2 The role of positional encodings
@@ H.3 ER vs DAG performance
Directo-DD RRWP 79 $\pm$ 3.7 100 $\pm$ 0.0 79 $\pm$ 3.7
Directo-DD MagLap ($Q$ = 10) 86 $\pm$ 9.4 99 $\pm$ 1.2 85 $\pm$ 10.2
Directo RRWP 94 $\pm$ 1.0 100 $\pm$ 0.0 94 $\pm$ 1.0
Directo MagLap ($Q$ = 10) 99.5 $\pm$ 1.0 92.5 $\pm$ 2.7 92 $\pm$ 3.7
@@
H.4 Effect of dataset size
@@
Table 14: Effect of the dataset size (using Directo-DD).
@@
Standard Train set 0.0113 0.0103 0.0355 0.0038 0.0024 1.0 99.2 100 0 0
Lap 0.0138 $\pm$ 0.0038 0.0144 $\pm$ 0.0042 0.0462 $\pm$ 0.0071 0.0055 $\pm$ 0.0011 0.0052 $\pm$ 0.0015 1.5 $\pm$ 0.3 67 $\pm$ 5.3 100 $\pm$ 0.0 100 $\pm$ 0.0 67 $\pm$ 5.3
RRWP 0.0142 $\pm$ 0.0040 0.0129 $\pm$ 0.0034 0.0408 $\pm$ 0.0053 0.0055 $\pm$ 0.0010 0.0048 $\pm$ 0.0014 1.4 $\pm$ 0.3 79 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 79 $\pm$ 3.7
MagLap 0.0142 $\pm$ 0.0044 0.0135 $\pm$ 0.0034 0.0438 $\pm$ 0.0071 0.0057 $\pm$ 0.0013 0.0046 $\pm$ 0.0010 1.4 $\pm$ 0.3 69 $\pm$ 6.4 100 $\pm$ 0.0 100 $\pm$ 0.0 69 $\pm$ 6.4
@@
H.5 Impact of sampling optimization using discrete flow matching

### v2

@@ H.3 The role of dual attention
Table 12: Directed Graph Generation performance across different transformer architectures using discrete flow matching. MagLap considers $Q=1$.
@@
RRWP - Dual 0.0107 $\pm$ 0.0019 0.0104 $\pm$ 0.0012 0.1209 $\pm$ 0.0194 0.0054 $\pm$ 0.0005 0.0043 $\pm$ 0.0008 1.7 $\pm$ 0.1 94.0 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 94.0 $\pm$ 1.0
MagLap 0.0113 $\pm$ 0.0024 0.0110 $\pm$ 0.0015 0.1196 $\pm$ 0.0107 0.0061 $\pm$ 0.0011 0.0029 $\pm$ 0.0006 1.7 $\pm$ 0.2 74.0 $\pm$ 4.4 100 $\pm$ 0.0 100 $\pm$ 0.0 74.0 $\pm$ 4.4
MagLap - Dual 0.0110 $\pm$ 0.0017 0.0116 $\pm$ 0.0021 0.0509 $\pm$ 0.0017 0.0062 $\pm$ 0.0014 0.0027 $\pm$ 0.0005 1.3 $\pm$ 0.2 91.0 $\pm$ 2.5 100 $\pm$ 0.0 100 $\pm$ 0.0 91.0 $\pm$ 2.5
SBM No PE 0.0038 $\pm$ 0.0021 0.0040 $\pm$ 0.0020 0.1912 $\pm$ 0.0178 0.0975 $\pm$ 0.0028 0.0961 $\pm$ 0.0071 20.5 $\pm$ 2.0 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
@@
MagLap - Dual 0.0039 $\pm$ 0.0011 0.0036 $\pm$ 0.0010 0.0653 $\pm$ 0.0026 0.0033 $\pm$ 0.0003 0.0038 $\pm$ 0.0010 1.9 $\pm$ 0.3 77.0 $\pm$ 7.6 100 $\pm$ 0.0 100 $\pm$ 0.0 77.0 $\pm$ 7.6
In particular, we observe that, across both datasets, the inclusion of the dual attention mechanism consistently improves the validity of the generated graphs and reduces structural discrepancies as measured by clustering, spectral, and wavelet distances. In the ER-DAG setting, models with dual attention (e.g., RRWP-Dual and MagLap-Dual) significantly outperform their non-dual counterparts in validity, achieving up to 94% validity with RRWP-Dual and 91% with MagLap-Dual, compared to just 67.5% and 74%, respectively. These models also demonstrate better structural fidelity, particularly in clustering and spectral metrics. A similar trend is observed in the SBM dataset, where dual attention boosts validity (e.g., RRWP-Dual: 87%, MagLap-Dual: 77%). These results demonstrate how incorporating bidirectional information flow into the attention mechanism contributes to improved model performance and provides insight into the effectiveness of this architectural component.
Directo-$D$
@@
Table 13: Directed Graph Generation performance across different transformer architectures using discrete diffusion. MagLap considers $Q=1$.
@@
ER-DAG No PE 0.0094 $\pm$ 0.0024 0.0095 $\pm$ 0.0023 0.1221 $\pm$ 0.0122 0.0814 $\pm$ 0.0030 0.1094 $\pm$ 0.0066 14.4 $\pm$ 0.8 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
No PE - Dual 0.0149 $\pm$ 0.0037 0.0138 $\pm$ 0.0035 0.1256 $\pm$ 0.0088 0.0089 $\pm$ 0.0018 0.0106 $\pm$ 0.0014 2.6 $\pm$ 0.3 53.0 $\pm$ 2.9 100 $\pm$ 0.0 100 $\pm$ 0.0 53.0 $\pm$ 2.9
RRWP 0.0147 $\pm$ 0.0038 0.0139 $\pm$ 0.0038 0.0985 $\pm$ 0.0154 0.0067 $\pm$ 0.0009 0.0038 $\pm$ 0.0009 2.1 $\pm$ 0.3 42.0 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 42.0 $\pm$ 1.0
RRWP - Dual 0.0142 $\pm$ 0.0040 0.0129 $\pm$ 0.0034 0.0408 $\pm$ 0.0053 0.0055 $\pm$ 0.0010 0.0048 $\pm$ 0.0014 1.4 $\pm$ 0.3 79.0 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 79.0 $\pm$ 3.7
MagLap 0.0144 $\pm$ 0.0040 0.0135 $\pm$ 0.0038 0.0792 $\pm$ 0.0126 0.0071 $\pm$ 0.0012 0.0036 $\pm$ 0.0011 1.6 $\pm$ 0.3 59.0 $\pm$ 6.8 100 $\pm$ 0.0 100 $\pm$ 0.0 59.0 $\pm$ 6.8
MagLap - Dual 0.0142 $\pm$ 0.0044 0.0135 $\pm$ 0.0034 0.0438 $\pm$ 0.0071 0.0057 $\pm$ 0.0013 0.0046 $\pm$ 0.0010 1.4 $\pm$ 0.3 69.0 $\pm$ 6.4 100 $\pm$ 0.0 100 $\pm$ 0.0 69.0 $\pm$ 6.4
SBM No PE 0.0038 $\pm$ 0.0021 0.0040 $\pm$ 0.0020 0.1912 $\pm$ 0.0177 0.0120 $\pm$ 0.0019 0.0975 $\pm$ 0.0081 20.5 $\pm$ 2.0 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
No PE - Dual 0.0037 $\pm$ 0.0020 0.0039 $\pm$ 0.0019 0.0721 $\pm$ 0.0059 0.0055 $\pm$ 0.0007 0.0403 $\pm$ 0.0067 8.8 $\pm$ 1.3 35.0 $\pm$ 9.1 100 $\pm$ 0.0 100 $\pm$ 0.0 35.0 $\pm$ 9.1
RRWP 0.0038 $\pm$ 0.0019 0.0038 $\pm$ 0.0018 0.1085 $\pm$ 0.0205 0.0088 $\pm$ 0.0009 0.0359 $\pm$ 0.0043 8.5 $\pm$ 0.3 4.0 $\pm$ 2.0 100 $\pm$ 0.0 100 $\pm$ 0.0 4.0 $\pm$ 2.0
@@
H.4 The role of positional encodings
@@ H.5 ER vs DAG performance
Directo-DD RRWP 79.0 $\pm$ 3.7 100 $\pm$ 0.0 79.0 $\pm$ 3.7
Directo-DD MagLap ($Q$ = 10) 86.0 $\pm$ 9.4 99 $\pm$ 1.2 85.0 $\pm$ 10.2
Directo RRWP 94.0 $\pm$ 1.0 100 $\pm$ 0.0 94.0 $\pm$ 1.0
Directo MagLap ($Q$ = 10) 99.5 $\pm$ 1.0 92.5 $\pm$ 2.7 92.0 $\pm$ 3.7
@@
H.6 Scalability experiments
In this section, we investigate the scalability of our method with respect to three key axes: dataset size, parameter efficiency, and graph size. These experiments are designed to disentangle the contributions of data availability, architectural choices, and input complexity to the overall generative performance, providing a clearer understanding of the trade-offs between accuracy, efficiency, and computational cost in our framework.
Effect of dataset size
@@ H.6 Scalability experiments
Table 17: Effect of the dataset size (using Directo-$D$).
@@
Standard Train set 0.0113 0.0103 0.0355 0.0038 0.0024 1.0 99.2 100 0.0 0.0
Lap 0.0138 $\pm$ 0.0038 0.0144 $\pm$ 0.0042 0.0462 $\pm$ 0.0071 0.0055 $\pm$ 0.0011 0.0052 $\pm$ 0.0015 1.5 $\pm$ 0.3 67.0 $\pm$ 5.3 100 $\pm$ 0.0 100 $\pm$ 0.0 67.0 $\pm$ 5.3
RRWP 0.0142 $\pm$ 0.0040 0.0129 $\pm$ 0.0034 0.0408 $\pm$ 0.0053 0.0055 $\pm$ 0.0010 0.0048 $\pm$ 0.0014 1.4 $\pm$ 0.3 79.0 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 79.0 $\pm$ 3.7
MagLap 0.0142 $\pm$ 0.0044 0.0135 $\pm$ 0.0034 0.0438 $\pm$ 0.0071 0.0057 $\pm$ 0.0013 0.0046 $\pm$ 0.0010 1.4 $\pm$ 0.3 69.0 $\pm$ 6.4 100 $\pm$ 0.0 100 $\pm$ 0.0 69.0 $\pm$ 6.4
@@
Effect of model size (parameter efficiency)
We next evaluate parameter efficiency by contrasting our dual-attention mechanism with an alternative approach that scales model capacity by doubling the number of parameters. Dual attention increases the parameter space by introducing two adjacency-based attention heads, each specialized for handling edge directionality. To ablate for model size, we compare against a baseline in which the architecture is widened to match the parameter count, but without dual-attention.
Table 18: Directed Graph Generation performance for the dual attention architecture versus doubling the depth of the network without using the dual attention mechanism.
ER-DAG RRWP - Double 0.0129 $\pm$ 0.0020 0.0130 $\pm$ 0.0015 0.1741 $\pm$ 0.0200 0.0107 $\pm$ 0.0007 0.0020 $\pm$ 0.0005 4.9 $\pm$ 0.2 72.0 $\pm$ 9.8 100 $\pm$ 0.0 100 $\pm$ 0.0 72.0 $\pm$ 9.8
RRWP - Dual 0.0107 $\pm$ 0.0019 0.0104 $\pm$ 0.0012 0.1209 $\pm$ 0.0194 0.0054 $\pm$ 0.0005 0.0043 $\pm$ 0.0008 1.7 $\pm$ 0.1 94.0 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 94.0 $\pm$ 1.0
MagLap - Double 0.0131 $\pm$ 0.0019 0.0118 $\pm$ 0.0022 0.1237 $\pm$ 0.0217 0.0095 $\pm$ 0.0030 0.0020 $\pm$ 0.0008 4.3 $\pm$ 0.7 80.0 $\pm$ 6.3 100 $\pm$ 0.0 100 $\pm$ 0.0 80.0 $\pm$ 6.3
MagLap - Dual 0.0110 $\pm$ 0.0017 0.0116 $\pm$ 0.0021 0.0509 $\pm$ 0.0017 0.0062 $\pm$ 0.0014 0.0027 $\pm$ 0.0005 1.3 $\pm$ 0.2 91.0 $\pm$ 2.5 100 $\pm$ 0.0 100 $\pm$ 0.0 91.0 $\pm$ 2.5
SBM RRWP - Double 0.0083 $\pm$ 0.0024 0.0081 $\pm$ 0.0025 0.1789 $\pm$ 0.0107 0.0084 $\pm$ 0.0014 0.0084 $\pm$ 0.0067 27.1 $\pm$ 2.3 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
RRWP - Dual 0.0038 $\pm$ 0.0011 0.0037 $\pm$ 0.0012 0.0492 $\pm$ 0.0041 0.0035 $\pm$ 0.0007 0.0015 $\pm$ 0.0004 1.4 $\pm$ 0.2 87.0 $\pm$ 5.1 100 $\pm$ 0.0 100 $\pm$ 0.0 87.0 $\pm$ 5.1
MagLap - Double 0.0081 $\pm$ 0.0025 0.0079 $\pm$ 0.0020 0.1799 $\pm$ 0.0097 0.0307 $\pm$ 0.0052 0.0723 $\pm$ 0.0156 24.8 $\pm$ 4.6 6.0 $\pm$ 4.9 100 $\pm$ 0.0 100 $\pm$ 0.0 6.0 $\pm$ 4.9
MagLap - Dual 0.0039 $\pm$ 0.0011 0.0036 $\pm$ 0.0010 0.0653 $\pm$ 0.0026 0.0033 $\pm$ 0.0003 0.0038 $\pm$ 0.0010 1.9 $\pm$ 0.3 77.0 $\pm$ 7.6 100 $\pm$ 0.0 100 $\pm$ 0.0 77.0 $\pm$ 7.6
As shown in Table 18, this baseline fails to recover the correct edge distribution, particularly in more challenging regimes such as SBM graphs. In contrast, the dual-attention model achieves substantially higher fidelity, demonstrating that our architecture leverages parameters more effectively by embedding structural inductive bias rather than merely increasing model scale.
Effect of node size
We finally investigate how the method scales with increasing graph size by conducting experiments on ER-DAG datasets ranging from 80–150 up to 200–250 nodes (see Table 19 for dataset statistics). Due to the poor scalability of MultMagLap encodings, we restrict this ablation to RRWP and MagLap positional encodings, with the largest graphs (200–250 nodes) evaluated only under RRWP. Training and sampling times, together with generation quality, are reported in Table 20. Importantly, the computational overhead introduced by dual attention remains bounded by a constant factor of two and does not affect the overall asymptotic complexity, which matches that of standard graph transformers.
Table 19: Dataset statistics for the synthetic experiments with larger graph size (nodes and edges).
Min. nodes Max. nodes Avg. nodes Min. edges Max. edges Avg. edges
80 150 117 1866 6699 4170
150 200 173 6643 11851 9046
200 250 225 11982 18645 15148
In terms of positional encodings, MagLap exhibits clear scalability limitations, both in runtime and performance, whereas RRWP features scale considerably better, consistent with prior findings (e.g., Appendix G.4 of DeFoG (Qin et al., 2025a)). Regarding generative performance, DIRECTO demonstrates strong distributional fidelity across node sizes, as reflected in the Ratio and distributional validity scores. However, ensuring strict acyclicity becomes increasingly challenging at scale: a single misplaced edge can violate the constraint entirely, and the risk grows quadratically with the number of nodes. This highlights the inherent difficulty of enforcing global constraints in large directed graphs, motivating future work on more robust inductive biases for acyclicity preservation.
Table 20: Scaling performance of Directo on ER-DAG datasets of increasing node size.
Model Out-degree $\downarrow$ In-degree $\downarrow$ Clustering $\downarrow$ Spectre $\downarrow$ Wavelet $\downarrow$ Ratio $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ Novel $\uparrow$ V.U.N. $\uparrow$ Train (h) Min / sample
80-150 (RRWP) 0.0045 $\pm$ 0.0001 0.0044 $\pm$ 0.0008 0.3233 $\pm$ 0.0676 0.0010 $\pm$ 0.0003 0.0003 $\pm$ 0.0001 3.7 $\pm$ 0.5 53.5 $\pm$ 3.0 100 $\pm$ 0.0 100 $\pm$ 0.0 53.5 $\pm$ 3.0 22 0.285
80-150 (MagLap) 0.0040 $\pm$ 0.0005 0.0044 $\pm$ 0.0005 0.3382 $\pm$ 0.0361 0.0015 $\pm$ 0.0002 0.0007 $\pm$ 0.0002 4.2 $\pm$ 0.3 45.0 $\pm$ 7.1 100 $\pm$ 0.0 100 $\pm$ 0.0 45.0 $\pm$ 7.1 26 0.940
150-200 (RRWP) 0.0034 $\pm$ 0.0002 0.0031 $\pm$ 0.0002 0.1570 $\pm$ 0.0496 0.003 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 1.7 $\pm$ 0.3 60.0 $\pm$ 5.9 100 $\pm$ 0.0 100 $\pm$ 0.0 60.0 $\pm$ 5.9 28 0.555
150-200 (MagLap) 0.0031 $\pm$ 0.0005 0.0031 $\pm$ 0.0004 0.6691 $\pm$ 0.0558 0.0004 $\pm$ 0.0000 0.0001 $\pm$ 0.0001 3.2 $\pm$ 0.4 34.5 $\pm$ 1.9 100 $\pm$ 0.0 100 $\pm$ 0.0 34.5 $\pm$ 1.9 38 1.365
200-250 (RRWP) 0.0034 $\pm$ 0.0002 0.0033 $\pm$ 0.0003 0.0411 $\pm$ 0.0178 0.0003 $\pm$ 0.0000 0.0000 $\pm$ 0.0000 1.9 $\pm$ 0.3 2.0 $\pm$ 2.4 100 $\pm$ 0.0 100 $\pm$ 0.0 2.0 $\pm$ 2.4 40 0.645
H.7 Conditional generation
Table 21 shows the results of the experiments with classifier-free guidance conditional generation. We followed the setup of Li et al. (2025), training on the TPU Tiles dataset with the provided conditional information and evaluating with their available suite and surrogate models.
Table 21: Averaged conditional generation results across 3 runs. Baseline results were obtained from Li et al. (2025), where Spearman correlation was not reported.
Model Pearson Spearman MAE
Real graphs 0.75 $\pm$ 0.01 0.83 $\pm$ 0.01 0.96 $\pm$ 0.00
D-VAE 0.50 $\pm$ 0.01 N/A 1.4 $\pm$ 0.00
GraphRNN 0.62 $\pm$ 0.02 N/A 1.3 $\pm$ 0.00
GraphPNAS 0.24 $\pm$ 0.10 N/A 2.1 $\pm$ 0.06
OneShotDAG 0.56 $\pm$ 0.02 N/A 1.4 $\pm$ 0.01
LayerDAG (T=1) 0.37 $\pm$ 0.11 N/A 2.0 $\pm$ 0.04
LayerDAG 0.65 $\pm$ 0.01 N/A 1.2 $\pm$ 0.01
Directo 0.59 $\pm$ 0.01 0.63 $\pm$ 0.03 1.3 $\pm$ 0.01
Unlike methods explicitly designed for conditional DAG generation, our framework is general and primarily optimized for unconditional digraph generation. In contrast, autoregressive models such as GraphRNN and LayerDAG (especially the latter, where conditional generation is the default training setup) are naturally more suited for this task and thus represent strong baselines. Moreover, our evaluation follows their official suite, which is tailored to their own method and may not fully align with our formulation.
Despite this, Directo achieves competitive performance. In Pearson correlation, DIRECTO ranks just behind LayerDAG and GraphRNN, while performing on par with them in mean absolute error (MAE). Notably, it also outperforms LayerDAG’s one-shot variant OneShotDAG. We further report Spearman correlation, which is implemented in the evaluation suite but was not included in prior work. These results show that even though conditional generation is not our main focus, DIRECTO achieves a strong balance across all metrics, demonstrating that our framework can readily extend to conditional tasks without specialized architectural modifications.
Finally, our results were obtained using classifier-free guidance, a straightforward extension that may not be optimal for conditional training in our setting. We expect that more dedicated strategies for conditioning could further improve performance. Overall, these findings highlight the versatility of DIRECTO: while unconditional generation remains our primary focus, its ability to adapt effectively to conditional generation underscores the promise of diffusion-based formulations for directed graph generation.
H.8 Impact of sampling optimization using discrete flow matching